What is the wavenumber of IR radiation designated by 13.04 microns?
Answer #1

13.04 microns  

                     = 13.04*0.0001cm               [ 1 micron = 0.0001cm]

                    = 0.001304cm

wave number is reciprocal of wave length

wavenumber   = 1/0.001304cm     =   766.87 cm^-1 >>>>answer